# Rotary Outliers and Rotary Offset Features in Large Language Models

### v1

@@ 4.3.2 Metrics across models
DeepSeek-V2-Lite and Llama-2-7b also have rotary offset features, similarly to Phi-1, despite their different architectures. Table 2 presents the recall metrics for both upper and lower bounds across our analyzed models. For the upper bound, we find that most outliers exist in the rotary offset features, with over 90% of the outlier features being in potential rotary offset features selected by the upper bound. The lower bound, despite covering only approximately 35% of possible angles within roughly 30% of all features, successfully captures the majority of outlier features across all three models. Visual evidence in Figures 8 (and Figure 11 for DeepSeek, in Appendix) further supports this finding, showing clear clustering of large radius features around our predicted lower bounds. This clustering provides strong evidence that our derived lower bounds accurately identify effective rotary offset features and that they tend to contain large magnitude features. Relaxing our lower bound by 0.1 rad or $\sim$6 we see that the recall for the lower bound increases, indicating that large amounts of outlier features are also just below the lower bound.
5 Discussion
Our predictions about the behaviour of rotary offset features and the empirical results from our bounds suggest that the main drivers for outlier features in low-frequency rotary features are rotary offset features. The rotary pairs of these features correspond to the two high-norm bands seen in Figure 1.
@@ 6 Related Work
(Chen et al., 2024; Liu et al., 2024a) studies length generalization in RoPE and hypothesizes that low-frequency features are unable to generalize to lengths beyond the context length. They derive an upper bound for the frequency of these features which turns out to be the same as the upper bound for our rotary offset features. We independently discovered this upper bound during our research on rotary outlier features and show how it can instead be applied to predict which rotary frequencies are likely to contain outlier activations.
@@
Recent work has identified attention sinks (Xiao et al., 2024) as a key mechanism in language models, where certain tokens (typically beginning of the sequence and semantically unimportant tokens) persistently receive high attention scores across many attention heads. Our analysis reveals the rotary offset features as one mechanism of how attention sinks are consistently selected for across positions. (Gu et al., 2024) also notes that attention sinks emerge when the cosine between queries and keys is positive and large. We show how this works on an individual query-key level and how it relates to the rotary offset features.
(Liu et al., 2024b) studies how to improve kv-cache quantization. They note that high-norm channels in keys are not present in the value vectors. They choose to employ asymmetric channel-wise quantization for keys, among other techniques, to reduce the impact of these high-norm channels on the model’s performance after quantization.
@@ 7 Conclusion
First, we demonstrate that certain rotary feature pairs consistently emerge as outliers with large radii across different models and architectures. These outliers serve specific positional attention functions - low frequency features enable rotary offset features while high frequency features create precise positional patterns like sub-diagonal attention.
Second, we derive bounds that explain why lower frequency rotary features can effectively implement rotary features that offset the attention score with distance and what angle their rotary pairs’ queries and keys tend to be above. This matches empirical observations of where outlier features emerge.
We hope the derived bounds, the insights and the analysis presented in this paper can help further the understanding of models using rotary positional embeddings, lead to improvements in the frequency selection of rotary features, and potentially inform the development of improved quantization algorithms for language models using rotary embeddings.

### v2

@@ 4.3.2 Metrics across models
DeepSeek-V2-Lite and Llama-3-8B/70B also have rotary offset features, similarly to Phi-1, despite their different architectures and scales. Table 2 presents the recall metrics for both upper and lower bounds across our analyzed models. For the upper bound, we find that features with radius greater than radius 6.0 are within the bounds over 90% of the time in Phi-1, DeepSeek-V2-Lite and Llama-3-8B. The lower bound, despite covering less than 40% of possible angles between 28-45% of all features, depending on the model, successfully captures the majority of large-radius features across Phi-1, DeepSeek-V2-Lite and Llama-3-8B. For Llama-3-70B, we see lower recall for radius 6.0 than for the other models, however from radius 9.0 and greater our bounds show good recall, and we again see a clustering around the lower bound since there’s almost 20% of features with large radius just below the lower bound. Visual evidence in Figures 8 (and Figure 11 for DeepSeek) further supports our findings, showing clear clustering of large-radius features around our predicted lower bounds. Relaxing our lower bound by 0.1 rad or $\sim$6 the recall for the lower bound improves, indicating that a concentration of large-radius features are also just below the lower bound.
Our predictions about the behaviour of rotary offset features and the empirical results from our bounds suggest that one of the main drivers for large-magnitude features in low-frequency rotary features are rotary offset features. The pair of feature dimensions that make up the rotary offset features correspond to the two high-norm bands seen in Figure 1.
@@ 5 Related Work
(Chen et al., 2024; Liu et al., 2024b) studies length generalization in RoPE and hypothesizes that low-frequency features are unable to generalize to lengths beyond the context length. They derive an upper bound for the frequency of these features which turns out to be the same as the upper bound for our rotary offset features. We independently discovered this upper bound during our research on large-radius rotary features and show how it can instead be applied to predict which rotary frequencies are likely to contain large-magnitude activations.
@@
Recent work has identified attention sinks (Xiao et al., 2024) as a key mechanism in language models, where certain tokens (typically beginning of the sequence and semantically unimportant tokens) persistently receive high attention scores across many attention heads. Our analysis reveals the rotary offset features as one mechanism of how attention sinks are consistently selected for across positions. (Gu et al., 2024) also notes that attention sinks emerge when the cosine between queries and keys is positive and large. We show how this works on an individual query-key level for the rotary offset features.
(Liu et al., 2024c) studies how to improve kv-cache quantization. They note that high-norm channels in keys are not present in the value vectors. They choose to employ asymmetric channel-wise quantization for keys, among other techniques, to reduce the impact of these high-norm channels on the model’s performance after quantization.
@@ 6 Conclusion
First, we demonstrate that certain rotary feature pairs consistently emerge with large radii across different models and architectures. These features tend to have a U-shaped contribution to the attention score with relative distance between tokens. We call these rotary feature pairs rotary offset features.
Second, we derive bounds that predict at which rotary frequencies rotary offset features emerge and above what angle the corresponding rotary feature pairs in the queries and keys tend to be. This matches empirical observations of where large-radius rotary feature pairs appear. These bounds solely rely on context length and rotary frequency as input.
We hope the derived bounds, the insights and the analysis presented in this paper can help further the understanding of models using rotary positional embeddings, lead to improvements in the frequency selection of rotary features for context length extension and inform the development of improved quantization algorithms for language models using rotary embeddings.
@@
Appendix A Additional Figures for main text
Figure 9: Query and key magnitudes prior to rotation in Phi-1. The red dashed line separates rotary (left) from non-rotary (right) features, we further separate the first dimensions of the rotary pairs (left) from the second dimensions (right) with a white dashed line. The bands of large absolute values are located at the pair of dimensions that are rotated by the same low frequency in a sliced rotary implementation layout.
Figure 10: Attention pattern after applying scaling, masking and softmax to $\Sigma_{i}d_{i}(p)$ shown as total in Fig 4.
Figure 11: DeepSeek-V2-Lite: Mean radius of keys plotted against the angle between mean query and mean key for each layer (colored) and head for the last 12 rotary features. Black dashed lines for feature 23-31 are the lower bounds for the angle $\phi_{i}$ for each feature to serve as a rotary offset feature. Any feature without a dashed line are outside the bounds. The rotational keys are repeated 16 times to match the queries, so there are 16 times as many datapoints as there are actual rotational keys. We do note a single feature, feature 21, which is a low-frequency cyclic rotary feature that has feature instances with large radius in some heads. It may be the result of context length extension (see Section C for a potential explanation).
Appendix B Comparing Query-Key distributions for Wikipedia and Code Sample
Figure 12: The x and y coordinates of the query and key vectors for features 0-15 in Phi-1 in the same head analyzed in the main text.
Figure 13: The x and y coordinates of the query and key vectors for features 0-15 in Phi-1 in the same head analyzed in the main text. But using the code sample instead of the wiki sample. Compare against Figure 12.
Figure 14: Mean radius of keys plotted against the angle between mean query and mean key for each layer (colored) and head for the last 12 rotary features using the code sample for Phi-1. Same plot as in Figure 8 but using the code sample for collecting activations.
The analysis in the main text uses a data sample from Wikipedia, however the properties hold for many other samples. In Figures 12 and 13 we collect the activations for the same data as the main text and also for a code sample and plot the queries and keys for the rotary dimensions of Phi-1. Here we can clearly see that Feature 12 maintains the distributions and the angle of the queries and the keys between two different data samples. In Figure 14 we reproduce the same plot as in Figure 8 but for the code sample to show that the properties hold for other data samples.
Appendix C Llama-3 series and context length extension
Figure 15: A comparison of the key mean radius and angle differences between the base and extended context length versions of Llama 8B 3 and 3.1 for features 39 and 45. In both cases, it can be observed that fine-tuning for extension moves the features closer to the extended bounds. Features with greater margin after extension (45) are seen to increase in radius and features with less margin (39) are seen to decrease in radius.
Figure 16: A comparison of the key mean radius for features that are no longer candidates for being rotary offset features after extension (features 35-39) and features that are candidates for being rotary offset features both before and after extension (features 40-63) for every head in the models. With extension fine-tuning the sum of the radii decreases for features 35-39 and increases for 40-63.
We compare the basic and extended context length versions of Llama 8B 3 and 3.1 respectively. We firstly find that the basic version follows our bounds. The extended version, at first glance, looks to largely follow the base model’s bounds instead of the extended bounds. Looking closer, we find that the features that after the extension have become outside or with low margin (close to $2\pi$) to our upper bound are decreasing in radius while the features that have an upper bound with larger margin are increasing in radius. The extended version’s angle differences are also shifted closer to the extended bounds. Figure 15 shows the key mean radius and angle differences between the base and extended context length versions of Llama 8B 3 and 3.1 for features 39 and 45. Figure 16 shows that the pattern seen in Figure 15 with decreasing radii for features 35-39 and increasing radii for 40-63 is a pattern that holds across many heads.
In summary, some context length extension methods produce models that may not follow their new bounds, but instead follow something inbetween the base and extended bounds. The evolution of the radii and angular differences suggest the context length extension fine-tuning process may be adjusting the model towards the new extended bounds, but it is unclear whether it would converge with enough data. DeepSeek-V2-Lite is also a context length extended (which uses YaRN extension (Peng et al., 2023)) model, where our bounds fit well with the exception of feature 21. It is unclear whether feature 21 is a result of their context length extension method.
Model dh r pmax extended
Llama-3-8B 128 128 4096 No
Llama-3.1-8B 128 128 131072 Yes
Table 3: Model architecture details showing head dimension ($d_{h}$), rotary dimension ($r$), context length ($p_{max}$), and whether the linear projections for keys and queries have bias and whether the model was extended for longer context length.
A potential future research direction for making LLMs adopt context length extension more effectively could be to alter the distribution of rotary frequencies. The altered distribution should make sure all the rotary frequencies that would be effective as rotary offset features at the original context length are also effective at the extended context length. This could be accomplished by, for instance, having high frequency rotary features replace the first few rotary offset features at the base context length (thus making them invalid as rotary offset features at both context lengths) up to the point where the extended version of the upper bound starts.
Appendix D Data used for collecting activations
We use the following text as input for collecting activation data from the models. It is an excerpt from a Wikipedia article on the United States of America. It is nearly 888 tokens if tokenized by Llama-3-8B tokenizer.
The United States of America (USA), also known as the United States (U.S.) or America, is a country primarily located in North America. It is a federal republic of 50 states and Washington, D.C. as its federal capital district. The 48 contiguous states border Canada to the north and Mexico to the south, with the semi-exclavic state of Alaska in the northwest and the archipelagic state of Hawaii in the Pacific Ocean. The U.S. also asserts federal sovereignty over five major island territories and various uninhabited islands.[k] In addition, some 326 Indian reservations are treated as ”domestic dependent nations” with tribal sovereignty rights. The U.S. is a megadiverse country, with the world’s third-largest land area[d] and third-largest population, exceeding 340 million.[l] Its three largest metropolitan areas are New York, Los Angeles, and Chicago, and its three most populous states are California, Texas, and Florida.
Paleo-Indians migrated to North America across the Bering land bridge more than 12,000 years ago, and formed various civilizations and societies. Spanish exploration and colonization led to the establishment in 1513 of Spanish Florida, the first European colony in what is now the continental United States. France also began to colonize at this time, but major settlements came much later. Subsequent British colonization led to the first settlement of the Thirteen Colonies in Virginia in 1607. Intensive agriculture in the rapidly expanding Southern Colonies encouraged the forced migration of enslaved Africans. Clashes with the British Crown over taxation and political representation sparked the American Revolution, with the Second Continental Congress formally declaring independence on July 4, 1776.
The United States emerged victorious from the 1775–1783 Revolutionary War and expanded westward across North America, dispossessing Native Americans as it fought the Indian Wars. Expansion continued when the U.S. signed the 1803 Louisiana Purchase with Napoleonic France and won the Mexican–American War in 1848. As more states were admitted, a North–South division over slavery led to the secession of the Confederate States of America, which fought the Union in the 1861–1865 American Civil War. With the victory and preservation of the United States, slavery was abolished nationally. By the late 19th century, the United States established itself as a great power with victory in the Spanish–American War, a status solidified with its participation in World War I. Following Japan’s attack on Pearl Harbor in December 1941, the U.S. entered World War II. The aftermath of the war left the U.S. and the Soviet Union as the world’s two superpowers which led to the Cold War, during which both countries struggled for ideological dominance and international influence. The end of the Cold War and the Soviet Union’s collapse in 1991 left the U.S. as the world’s sole superpower, with significant geopolitical influence globally.
The U.S. national government is a presidential constitutional federal republic and liberal democracy with three separate branches: legislative, executive, and judicial. It has a bicameral national legislature composed of the House of Representatives, a lower house based on population, and the Senate, an upper house based on equal representation for each state. The country’s Democratic and Republican parties have dominated American politics since the 1850s. Federalism provides substantial autonomy to the 50 states, while American values are based on a political tradition that draws its inspiration from the European Enlightenment movement. A melting pot of many ethnicities and customs, the culture of the United States has been shaped by centuries of immigration, and its soft power influence has a global reach.
One of the world’s most developed countries, the U.S. ranks among the highest in economic competitiveness, productivity, innovation, human rights, and higher education. The United States accounted for over a quarter of nominal global economic output in 2024, and its economy has been the world’s largest by nominal GDP since about 1890. It possesses by far the largest amount of wealth of any country and has the highest disposable household income per capita among OECD countries, though U.S. wealth inequality is higher than in most other developed countries. The U.S. is a member of multiple international organizations and plays a leading role in global political, cultural, economic, and military affairs.
Source: Wikipedia article on the United States, accessed March 2025
We use the following text as input for collecting activation data from the models for the Code Sample (used for Figure 13). It is code generated by GPT-5.
Compact Python utility for parallel URL fetching with optional file caching.
from __future__ import annotations
import hashlib, sys, threading, urllib.request
from concurrent.futures import ThreadPoolExecutor, as_completed
from dataclasses import dataclass
from datetime import datetime, timezone
from pathlib import Path
from typing import List, Optional
@dataclass
class FetchResult:
    url: str; status: int; content: bytes; fetched_at: datetime
class FileCache:
    def __init__(self, directory: Path):
        self.dir = directory; self.dir.mkdir(parents=True, exist_ok=True)
    def _path(self, key: str) -> Path:
        return self.dir / f"{hashlib.sha256(key.encode()).hexdigest()}.cache"
    def get(self, key: str) -> Optional[bytes]:
        p = self._path(key); return p.read_bytes() if p.exists() else None
    def set(self, key: str, data: bytes) -> None:
        p = self._path(key); p.write_bytes(data)
def fetch_url(url: str, timeout: int = 10) -> FetchResult:
    req = urllib.request.Request(url, headers={"User-Agent": "pyfetch/1.0"})
    try:
        with urllib.request.urlopen(req, timeout=timeout) as r:
            return FetchResult(url, getattr(r, "status", 200), r.read(), datetime.now(timezone.utc))
    except Exception as e:
        return FetchResult(url, 599, str(e).encode(), datetime.now(timezone.utc))
def parallel_fetch(urls: List[str], workers: int = 6, cache: Optional[FileCache] = None) -> List[FetchResult]:
    results = []
    with ThreadPoolExecutor(max_workers=workers) as ex:
        futures = {ex.submit(fetch_url, u): u for u in urls if not (cache and cache.get(u) and results.append(FetchResult(u,200,cache.get(u),datetime.now(timezone.utc))))}
        for fut in as_completed(futures):
            res = fut.result(); results.append(res)
            if cache and res.status == 200: cache.set(res.url, res.content)
    return results
def main(argv=None) -> int:
    argv = argv or sys.argv[1:]
    if not argv: return 2
    cache_dir, urls = None, []
    if "--cache" in argv:
        idx = argv.index("--cache"); cache_dir = Path(argv[idx+1]); del argv[idx:idx+2]
    urls = argv; cache = FileCache(cache_dir) if cache_dir else None
    res = parallel_fetch(urls, cache=cache)
    for r in res:
        print(r.url, r.status, len(r.content))
    return 0
if __name__ == "__main__":
    raise SystemExit(main())
Source: GPT-5 generated code sample, accessed August 2025